OpenStudy (anonymous):

The roots of the equation \(x^3-7x^2+Cx -8=0\) form a G.P. Find all roots.

OpenStudy (anonymous):

Well how do i start?

OpenStudy (anonymous):

what do u mean by ..........form a G.P. ?

OpenStudy (anonymous):

the roots form a Geometric Progression

OpenStudy (anonymous):

u mean the roots are somthing like \(\alpha\) , \(\alpha \beta\) and \(\alpha\beta^2\) ?

OpenStudy (anonymous):

I guess so; i can't remember how to do this.

mathslover (mathslover):

Hmn so we have common ratio as \(\large{\beta}\) right?

mathslover (mathslover):

*just for example*

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

can the roots be something like \((\alpha +\beta)(\beta+\gamma)(\gamma+\alpha)\)?

mathslover (mathslover):

http://www.wolframalpha.com/input/?i=x3%E2%88%927x2%2BCx%E2%88%928%3D0%20&t=crmtb01 "Seems somewhat 'difficult'"

mathslover (mathslover):

C is constant right? @Omniscience

OpenStudy (anonymous):

well the \(C\) has to be solved afterwards; i assume its a constant

OpenStudy (anonymous):

u can do some thing like this \[(x-\alpha)(x-\alpha \beta)(x-\alpha \beta^2)=x^3-3x^2+Cx-8\] simplify the Left hand side and then compare the coefficients

mathslover (mathslover):

ok ... so seems that the given polynomial is cubic ..

OpenStudy (anonymous):

u got 3 variables.... a , r , C u got to form three equations assume roots to be a/r , a , ar

mathslover (mathslover):

@A.Avinash_Goutham I do think that C is ..constant?

OpenStudy (anonymous):

-.- then?

OpenStudy (anonymous):

well its not a complex; and im not sure how to find the roots..haven't done these in a long time

OpenStudy (chaise):

http://www.wolframalpha.com/input/?i=a%5E3-bx%5E2%2BCx-d%3D0 You have 3 solutions for x, label these: x1, x2, x3 x2/x1=x3/x2 for it to be geometric progression. Plug in your values and divide, solve simultaneously. C cannot be 0, and B cannot be 0.

OpenStudy (chaise):

Once you find x1 and x2. Divide them, then x1*(x2/x1), x2*(x3/x2)^2 and so fourth. I think this is right, I hope this is right :s

OpenStudy (anonymous):

well i dont think that its that easy..

OpenStudy (anonymous):

any given information about C (integer or...)

OpenStudy (anonymous):

nope; i have to find \(C\) after that

OpenStudy (rsadhvika):

\(\alpha\beta = -2\)

OpenStudy (chaise):

I think I'm wrong. It won't be two equations with two unknown, because the values of x will be different.

OpenStudy (anonymous):

we know that that the product of roots for cubic polynomial equation like \(ax^3+bx^2+cx+d=0 \) is -d/a

OpenStudy (anonymous):

so according to the roots i assumed for ur equation \(\alpha^3 \beta^3=-8 \) so \(\alpha \beta=-2 \)

OpenStudy (anonymous):

sorry -d/a=8 \(\alpha \beta=2 \)

OpenStudy (anonymous):

wait how did you get \(2\)?

OpenStudy (anonymous):

product of roots=-d/a=8 am i right?

OpenStudy (anonymous):

i guess so..

OpenStudy (anonymous):

well the roots are \(\alpha\) , \(\alpha \beta\) and \(\alpha \beta^2\) and product of this 3 term is \(\alpha^3 \beta^3\)

OpenStudy (anonymous):

so \(\alpha \beta=2\)

OpenStudy (anonymous):

\(\alpha^3 \beta^3= (\alpha \beta)^3=8 \) then \(\alpha \beta=2\)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw :)